# Tool Unlearning for Tool-Augmented LLMs

### v1

@@ 6 Related work
Unlearning for non-LLM models: These methods include methods that focus on pruning before unlearning (Jia et al., 2023) or finding salient parameters (Fan et al., 2024b) and manipulating gradients Ullah et al. (2021); Hoang et al. (2024), adversarial methods (Liu et al., 2023; Setlur et al., 2022; Wei et al., 2023), approximation of inverse Hessian (Zhang et al., 2024a), and data augmentation (Choi et al., 2024). Other works study unlearning under multimodal setting (Cheng & Amiri, 2025), image-to-image models (Li et al., 2024a), and finding the most challenging unlearning subset within a dataset (Fan et al., 2024a). Recently, a few works started to benchmark MU performances on unlearning fictitious user profiles (Maini et al., 2024), world knowledge (Jin et al., 2024) and a variety of tasks (Cheng & Amiri, 2024).
@@
Tool-Augmented LLMs: TAML (Parisi et al., 2022) used self-play to boost LLMs’ performance on math and reasoning tasks. Schick et al. (2023) discovered that LLMs can teach themselves how to use APIs. Recently, efforts have been devoted to building benchmarks to train and evaluate the tool-using ability of LLMs, such as agent-based data generation (Tang et al., 2023; Li et al., 2023), bootstrapping training data with seed examples (Patil et al., 2023), modifying existing datasets (Basu et al., 2024), and dataset development with GPT-4 (Qin et al., 2024).
@@ 7 Conclusion
We introduce Tool Unlearning–a novel machine unlearning task with the goal of unlearning previously learned tools from tool-augmented LLMs. We develop the first tool unlearning approach, ToolDelete, that implements three key properties: tool knowledge deletion, tool knowledge retention, general capability retention. In addition, we introduce LiRA-Tool, the first membership inference attack (MIA) method for evaluating tool unlearning. LiRA-Tool largely addresses the limitations of sample-based MIAs for tool unlearning. Extensive experiments on several diverse datasets and LLMs show that ToolDelete is an efficient, flexible, and effective tool unlearning method that supports sequential unlearning, maintains strong performance across all key properties, and operates without requiring full access to training data. It outperforms existing methods by removing tool knowledge without over-forgetting (as shown in ablation studies), achieving 74.8% faster training times compared to retraining, and delivering highly effective tool unlearning even in resource-constrained settings with ToolDelete-LoRA (which reduces compute costs by 81.1% and training time by 71.3%).
In future, we will investigate tool unlearning in dynamically updated LLMs (e.g. API-based LLMs like GPT-4), where we address continuous unlearning challenges. In addition, we will develop adversarial training techniques and robustness evaluation frameworks to prevent unintended tool re-learning or model exploitation.
@@
We did not conduct experiments using closed-source LLMs or API-based LLMs. In addition, this work did not investigate the impact of varying model scales due to the limited publicly-available tool-augmented LLMs. Our experiments were conducted on the 7B scale and the scalability of the proposed tool unlearning approach across models of different sizes and scales is an open question for future investigation.
@@
Our work investigates the security implications of tool-augmented Large Language Models (LLMs), where we focus on the risks that arise from integrating external tools, and the necessity ability to remove these acquired tools. A key concern is ensuring compliance with privacy regulations, such as the Right to be Forgotten (RTBF), which mandates the removal of specific data upon user request. In the context of tool-augmented LLMs, this necessitates the ability to delete sensitive, regulated, or outdated information related to specific tools. By examining how LLMs interact with and rely on external tools, potential threats to model security can be identified, e.g. unauthorized tool usage, adversarial exploitation, and privacy violations. Our research highlights the critical importance of addressing these challenges.
@@ Appendix B Baselines
  • SalUn (Fan et al., 2024b) performs RandLabel on unlearning-related parameters discovered by saliency map.
@@
  • SGA (Jang et al., 2023; Barbulescu & Triantafillou, 2024), which performs gradient ascent on $\mathcal{T}_{f}$ whose memorization probability exceeds a pre-defined threshold.
  • TAU (Barbulescu & Triantafillou, 2024), which performs task arithmetic on SGA.
  • CUT (Li et al., 2024b).
@@ Appendix D Additional results
Pre-trained LLM weights $f_{0}$
ToolDelete-SFT 52.7 72.1 30.5 23.6
ToolDelete-DPO 53.4 75.1 28.7 23.1
Randomly initialized LLM $f_{R}$
ToolDelete-SFT 50.9 71.3 29.8 22.7
ToolDelete-DPO 52.6 73.4 27.5 22.4
@@ Appendix E Sampling of Shadow Samples for LiRA-Tool
You are now a synthetic data generator. Generate query-response pairs to evaluate an LLM’s ability of using an API. How to generate "query": Based on the API and documentation shown below, think of a user query that needs to be answered by calling the API. How to generate "response": Write down the correct API call with correct arguments. The in-context examples below demonstrate what you need to generate. Please be as diverse and creative as possible in phrasing and style. But do not hallucinate. ## In-context Examples #### Tool and Documentation Name: StableDiffusionPipeline.from_pretrained() #### Query I want to see some cats dancing in celebration! #### Response API call: StableDiffusionPipelin e.from_pretrained("stabilityai/stable-diffusion-2-1") Now, for the following API, generate a query-response pair. #### Tool and Documentation api_name() #### Query #### Response

### v2

@@ 6 Related work
Unlearning for non-LLM models: These methods include methods that focus on pruning before unlearning (Jia et al., 2023) or finding salient parameters (Fan et al., 2024) and manipulating gradients (Ullah et al., 2021; Hoang et al., 2024), adversarial methods (Liu et al., 2023; Setlur et al., 2022; Wei et al., 2023), approximation of inverse Hessian (Zhang et al., 2024a), and data augmentation (Choi et al., 2024). Other works study unlearning under multimodal setting (Cheng & Amiri, 2024a), image-to-image models (Li et al., 2024a), and finding the most challenging unlearning subset within a dataset (Fan et al., 2025b). Recently, a few works started to benchmark MU performances on unlearning fictitious user profiles (Maini et al., 2024), world knowledge (Jin et al., 2024) and a variety of tasks (Cheng & Amiri, 2024b).
@@
Tool-Augmented LLMs: Tool augmented language models (TAML) (Parisi et al., 2022) used self-play to boost LLMs’ performance on math and reasoning tasks. In addition, Toolformer (Schick et al., 2023) showed that LLMs can teach themselves how to use APIs. More recent efforts have been devoted to building benchmarks to train and evaluate the tool-using ability of LLMs. These include agent-based data generation (Tang et al., 2023; Li et al., 2023), bootstrapping training data with various seed examples (Patil et al., 2023), modifying existing datasets (Basu et al., 2024), and dataset development with powerfull LLMs such as GPT-4 (Qin et al., 2024).
@@ 7 Conclusion
We introduce Tool Unlearning–a novel machine unlearning task with the goal of unlearning previously learned tools from tool-augmented LLMs. We develop the first tool unlearning approach, ToolDelete, that implements three key properties: tool knowledge deletion, tool knowledge retention, general capability retention. In addition, we introduce LiRA-Tool, the first membership inference attack (MIA) method for evaluating tool unlearning. LiRA-Tool largely addresses the limitations of sample-based MIAs for tool unlearning. Extensive experiments on several diverse datasets and LLMs show that ToolDelete is an efficient, flexible, and effective tool unlearning method that supports sequential unlearning, maintains strong performance across all key properties, and operates without requiring full access to training data. It outperforms existing methods by removing tool knowledge without over-forgetting (as shown in ablation studies), achieving 74.8% faster training times compared to retraining, and delivering highly effective tool unlearning even in resource-constrained settings with ToolDelete-LoRA (which reduces compute costs by 81.1% and training time by 71.3%). In future, we will investigate tool unlearning in continually updated LLMs to address continuous unlearning challenges. In addition, we will develop adversarial training techniques and robustness evaluation frameworks to prevent unintended tool re-learning or model exploitation (Fan et al., 2025a), and conduct loss landscape analysis of tool unlearning (Cheng & Amiri, 2025)
@@
We did not conduct experiments using closed-source LLMs or API-based LLMs. In addition, this work did not investigate the impact of varying model scales due to the limited publicly-available tool-augmented LLMs. Our experiments were conducted on the 7B scale and the scalability of the proposed tool unlearning approach across models of different sizes and scales is an open question for future investigation. Moreover, evaluation of the efficacy of tool unlearning can be extended to broader conditions, such as under adversarial conditions (Łucki et al., 2025).
@@
Our work investigates machine unlearning in the context of tool-augmented Large Language Models (LLMs), where we focus on the risks that arise from integrating external tools and the crucial need for unlearning tool-usage capabilities for specific tools to ensure compliance with privacy regulations such as the Right to be Forgotten (RTBF). This necessitates the ability to delete sensitive, regulated, or outdated knowledge related to specific tools. Tool unlearning will enable us to identify potential threats to model security, e.g. unauthorized tool usage, adversarial exploitation, and privacy violations. Our research highlights the importance of addressing these challenges.
@@ Appendix B Baselines
  • SalUn (Fan et al., 2024) performs RandLabel on unlearning-related parameters discovered by saliency map.
@@
  • SGA (Jang et al., 2023; Bărbulescu & Triantafillou, 2024), which performs gradient ascent on $\mathcal{T}_{f}$ whose memorization probability exceeds a pre-defined threshold.
  • TAU (Bărbulescu & Triantafillou, 2024), which performs task arithmetic on SGA.
  • CUT (Li et al., 2024b), which controls model activations to be similar to the absence of knowledge on forget set.
@@ Appendix D Additional results
Pre-trained LLM weights $f_{0}$
ToolDelete-SFT 52.7 72.1 30.5 23.6
ToolDelete-DPO 53.4 75.1 28.7 23.1
Randomly initialized LLM $f_{R}$
ToolDelete-SFT 50.9 71.3 29.8 22.7
ToolDelete-DPO 52.6 73.4 27.5 22.4
@@ Appendix E Sampling of Shadow Samples for LiRA-Tool
You are now a synthetic data generator. Generate query-response pairs to evaluate an LLM’s ability of using an API. How to generate ”query”: Based on the API and documentation shown below, think of a user query that needs to be answered by calling the API. How to generate ”response”: Write down the correct API call with correct arguments. The in-context examples below demonstrate what you need to generate. Please be as diverse and creative as possible in phrasing and style. But do not hallucinate. ## In-context Examples #### Tool and Documentation Name: StableDiffusionPipeline.from_pretrained() #### Query I want to see some cats dancing in celebration! #### Response API call: StableDiffusionPipelin e.from_pretrained(”stabilityai/stable-diffusion-2-1”) Now, for the following API, generate a query-response pair. #### Tool and Documentation api_name() #### Query #### Response